OpenStudy (anonymous):

Tom receives allowance every week. He spends three-fourths of his allowance on food. Last week, Tom spent nine dollars on food. What is Tom’s weekly allowance? Write the verbal model, define the variable, write the equation and find the solution. Show your work.

OpenStudy (anonymous):

can someone please help

OpenStudy (anonymous):

9/3=3 so his weekly allowance would be twelve dollars since 12/4=3

OpenStudy (whpalmer4):

Let Tom's allowance be A. 3/4 of A is spent on food. Last week, he spent $9 on food, so \[9 = \frac{3}{4}A\] Solve that for \(A\).

OpenStudy (anonymous):

Convert 3/4 to a decimal . 100 divided by four equals 25. 25 times three equals 75. So your decimal is .75. Get a calculator and divide nine by .75 and see what you get. That will be A.

OpenStudy (whpalmer4):

That's kind of a lengthy way to get to the answer, @emilyowl! \[9 = \frac{3}{4}A\]Multiply both sides by 4:\[36 = 3A\]Divide both sides by 3:\[A =\]Shouldn't need a calculator...

OpenStudy (anonymous):

well i have the answer so never mind

OpenStudy (whpalmer4):

@312856MLP ah, that's the attitude that will certainly lead to long-term success!